OpenStudy (anonymous):

Solve the equation. If it is an identity (all real numbers) or a contradiction (no solution), so state. 3x + 5 + 6x = 9x + 8

OpenStudy (anonymous):

3x+5+6x=9x+8 9x+5=9x+8 9x=9x+5 0=5 No solution